OpenStudy (anonymous):

I've completed most of this question but I'm stuck, help please! Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+rn)1c]cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+rn.

OpenStudy (anonymous):

These are the first 2 questions and the answers I got. You are going to purchase a new car, but being a responsible consumer means doing a little bit of research first. First, you find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% 1. Using the function A(t)=P(1+rn)^nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. A(t) = 39,145(1 + 0.03/12)^12t 2. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal t = -45.54648631

OpenStudy (anonymous):

The last question is?

OpenStudy (anonymous):

The last one is the one I posted in my first post

OpenStudy (anonymous):

Dont you just take your original equation and put 4 whereever there is c?

OpenStudy (anonymous):

Here's how far I've gotten on the 3rd question: \[A(t)=39,145[(1+\frac{ .03 }{ 12 })^1/4]^(4)12t\] \[10,000=39,145[(1.0025)^1/4]^48t\]

OpenStudy (anonymous):

some of the variables are messed up...woops. (1.0025)^1/4]^48t (fixed for the second part)

OpenStudy (anonymous):

So where do you need help then

OpenStudy (anonymous):

I'm just confused as to what I do with the 1/4, I know that I would divide 39,145 like I did in problem #2 but after that I'm utterly confused

OpenStudy (anonymous):

is it 1/4 or ^1)/4

OpenStudy (anonymous):

I've done this question already, and there is a chance I still have it on my computer at home if you still have issues

OpenStudy (anonymous):

It's ^1/4 sorry for the late response my laptop died @swagmaster47

OpenStudy (anonymous):

Do you still need help or do you got it?

OpenStudy (anonymous):

@swagmaster I still need help lol

OpenStudy (anonymous):

Ok if you still need help this afternoon I will see what I did for that

OpenStudy (anonymous):

@swagmaster47 I think I figured it out. I believe the 1/4 reduces the 48t down to 12t which would basically make the answer the same as my previous question and the answer would be -45.54648631

OpenStudy (anonymous):

wait now o-o ugh I was thinking 1/2 for some reason. That's wrong.

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

oh wait no! I thought it divided but it multiplies to get 12, so yeah I was right! I've got it! @swagmaster47 thanks anyways :D I will give you a medal for being the only one who actually tried to help

OpenStudy (anonymous):

Oh thanks! I forgot about you sorry. I remember that problem being strange, thank you :)